A car travels at a speed of 44ft per second.what is this speed in miles per hour.
Found 2 solutions by nahataba, oberobic:
Answer by nahataba(3)

you find the conversion factor.
you set it up like this:
44ft/second and since there are 3600 seconds in an hour (60 minutes x 60 seconds) you times 44ft/1 second x 3600 seconds/ hour
you can then cross cancel the seconds (as in the words)
then since its miles per hour you would do 1 mi/5280 ft because thats how many feet are in a mile and you cross cancel the feet
so far this is what you should have:
44ft/1 x 3600/1 hr x 1 mi/5280
then you multiply 44 and 3600 which is 158,400 and you divide that by 5280 which is 30.
the answer is 30mi/hr

Answer by oberobic(2304)

44 ft/second = x mile/hr ??
.
44 ft = how many miles?
.
44 ft/sec / 5280 ft/mile = 1/120 mile/sec
.
1/120 mile/sec * 3600 sec/hr = 3600/120 mile/hr = 30 miles/hr
.
Done.